Sequence function(n)=n/n+1


1
Expert's answer
2021-05-26T15:38:14-0400

Prove that a sequence function(n)=n/n+1 converges to 1


Solution

"\lim\limits_{x\to \infin}\frac{n}{n+1}=1"


We need to show that for any given "\in \space \gt0,"  there is a natural number N such that if "n\gt N", then

"|\frac{n}{n+1}-1|=|\frac{-1}{n+1}|\lt \in"


NB: "\frac{1}{n+1}\lt\frac{1}{n}" for any "N\in \N"


for any given "\in" by Archimedean property, there exists "N\in \N" such that "\frac{1}{N}\lt\in."

Therefore, if "n\gt N" then, "\frac{1}{n+1}\lt\frac{1}{n}\lt\frac{1}{N}\lt \in"